# Are We Truly Forgetting? A Critical Re-examination of Machine Unlearning Evaluation Protocols

### v2

@@ A.1 Unlearning Baselines
We divide unlearning baselines into two categories based on whether they use retain set ($\mathcal{D}_{r}$) or not: (i) baselines without $\mathcal{D}_{r}$, and (ii) baselines with $\mathcal{D}_{r}$. The details of these algorithms are as follows.
@@
Unlearning w/o $\mathcal{D}_{r}$. In this section, baselines aim to remove information using only $\mathcal{D}_{f}$ without referencing retain set. While this reduces time complexity and the privacy risk, it often sacrifices some accuracy on retain set.
@@
Random Labeling (RL) [13]. RL discards ground-truth labels for forget set and replaces each $y_{i}$ with a label $y_{i}^{\mathrm{rand}}$ sampled from other classes:
@@
Pseudo Labeling (PL) [7]. PL ensures that $\mathcal{D}_{f}$ is unlearned in a manner that minimizes noise. For each sample $x_{i}\in\mathcal{D}_{f}$, the original model $\theta_{o}$ assigns a pseudo-label $y_{i}^{\prime}$, representing the class with the highest predicted probability after excluding forget set. The unlearning objective is defined as:
@@
Unlearning w/ $\mathcal{D}_{r}$. Several baselines utilize $\mathcal{D}_{r}$ as using only $\mathcal{D}_{f}$ often fails to preserve the knowledge of $\mathcal{D}_{r}$. Baselines in this category incorporate training or distillation on the retain set, thereby improving utility for the retain set.
@@
$\min_{\theta}$ $\sum_{(x_{i},y_{i})\in\mathcal{D}_{f}}\mathrm{dist}\Bigl{(}f_{\theta}(x_{i}),\mathbf{m}_{c^{*}}\Bigr{)},$ (6)
$c^{*}=$ $\arg\min_{c\neq y_{i}}\mathrm{dist}\bigl{(}f_{\theta}(x_{i}),\mathbf{m}_{c}\bigr{)}.$
@@
Contrastive Unlearning (CU) [38]. CU uses a contrastive objective to decouple the forget set from the retain set in embedding space. Though one can implement CU without $\mathcal{D}_{r}$ explicitly in every step (e.g., by approximating distances), it is often more effective when $\mathcal{D}_{r}$ is used to preserve the retain set.
@@
$\max_{\theta_{u}}$ $\sum_{(x_{j},y_{j})\in\mathcal{D}_{f}}d\bigl{(}f_{\theta_{u}}(x_{j}),f_{\theta}(x_{j})\bigr{)}$ (7)
$\quad+\min_{\theta_{u}}\;\sum_{(x_{j},y_{j})\in\mathcal{D}_{r}}d\bigl{(}f_{\theta_{u}}(x_{j}),f_{\theta}(x_{j})\bigr{)}.$
@@
SCAR [2]. SCAR exploits a distance-based realignment for $\mathcal{D}_{f}$, pushing their features toward incorrect classes to erase the original boundary, which preserves the representation for $\mathcal{D}_{r}$ via distillation or cross-entropy minimizing these terms:
$\sum_{(x_{i},y_{i})\in\mathcal{D}_{f}}\mathrm{dist}\bigl{(}f_{\theta}(x_{i}),\mathbf{m}_{c^{*}}\bigr{)}$ (8)
$\quad+\sum_{(x_{j},y_{j})\in\mathcal{D}_{r}}\ell\bigl{(}f_{\theta}(x_{j}),y_{j}\bigr{)}.$

### v3

@@ 3.5 Unlearning Baselines
We divide unlearning baselines into two categories based on whether they use the retain set ($\mathcal{D}_{r}$) or not: (i) baselines without $\mathcal{D}_{r}$, and (ii) baselines with $\mathcal{D}_{r}$. The details of these algorithms are as follows.
@@
Unlearning w/o $\mathcal{D}_{r}$. In this section, baselines aim to remove information using only $\mathcal{D}_{f}$ without the retain set. While this reduces time complexity and the privacy risk, it often sacrifices the accuracy on the retain set.
@@
Random Labeling (RL) [18]. RL discards ground-truth labels of the forget set and replaces each $y_{i}$ with a label $y_{i}^{\mathrm{rand}}$ sampled from other classes:
@@
Pseudo Labeling (PL) [8]. PL ensures that $\mathcal{D}_{f}$ is unlearned in a manner that minimizes noise. For each sample $x_{i}\in\mathcal{D}_{f}$, the original model $\theta_{o}$ assigns a pseudo-label $y_{i}^{\prime}$, representing the class with the highest predicted probability after excluding the forget set. The unlearning objective is defined as:
@@
Unlearning w/ $\mathcal{D}_{r}$. Several baselines utilize $\mathcal{D}_{r}$, as using only $\mathcal{D}_{f}$ often fails to preserve the knowledge of $\mathcal{D}_{r}$. Baselines in this category incorporate training or distillation on the retain set, thereby improving utility for the retain set.
@@
$\min_{\theta}$ $\sum_{(x_{i},y_{i})\in\mathcal{D}_{f}}\mathrm{dist}\Bigl(f_{\theta}(x_{i}),\mathbf{m}_{c^{*}}\Bigr),\ \ c^{*}=\;\arg\min_{c\neq y_{i}}\mathrm{dist}\bigl(f_{\theta}(x_{i}),\mathbf{m}_{c}\bigr).$ (8)
@@
Contrastive Unlearning (CU) [52]. CU uses a contrastive objective to decouple the forget set from the retain set in the embedding space. Though one can implement CU without $\mathcal{D}_{r}$ explicitly in every step (e.g., by approximating distances), it is often more effective when $\mathcal{D}_{r}$ is used to preserve the retain set.
@@
$\max_{\theta_{u}}$ $\sum_{(x_{j},y_{j})\in\mathcal{D}_{f}}d\bigl(f_{\theta_{u}}(x_{j}),f_{\theta}(x_{j})\bigr)+\min_{\theta_{u}}\sum_{(x_{j},y_{j})\in\mathcal{D}_{r}}d\bigl(f_{\theta_{u}}(x_{j}),f_{\theta}(x_{j})\bigr).$ (9)
@@
SCAR [3]. SCAR exploits a distance-based realignment for $\mathcal{D}_{f}$, pushing their features toward incorrect classes to erase the original boundary, which preserves the representation for $\mathcal{D}_{r}$ via distillation or cross-entropy, minimizing these terms:
$\sum_{(x_{i},y_{i})\in\mathcal{D}_{f}}\mathrm{dist}\bigl(f_{\theta}(x_{i}),\mathbf{m}_{c^{*}}\bigr)\quad+\sum_{(x_{j},y_{j})\in\mathcal{D}_{r}}\ell\bigl(f_{\theta}(x_{j}),y_{j}\bigr).$ (10)